### v2

Abstract
This paper presents Words2Contact, a language-guided multi-contact placement pipeline leveraging large language models and vision language models. Our method is a key component for language-assisted teleoperation and human-robot cooperation, where human operators can instruct the robots where to place their support contacts before whole-body reaching or manipulation using natural language. Words2Contact transforms the verbal instructions of a human operator into contact placement predictions; it also deals with iterative corrections, until the human is satisfied with the contact location identified in the robot’s field of view. We benchmark state-of-the-art LLMs and VLMs for size and performance in contact prediction. We demonstrate the effectiveness of the iterative correction process, showing that users, even naive, quickly learn how to instruct the system to obtain accurate locations. Finally, we validate Words2Contact in real-world experiments with the Talos humanoid robot, instructed by human operators to place support contacts on different locations and surfaces to avoid falling when reaching for distant objects.
I Introduction
Humanoid robots can use various body parts to create support contacts to help balance when reaching for difficult positions. For example, they can use their right hand as a support on a table, bend forward and reach a cup that would otherwise be out of reach (Fig. 1); or they can lean on the counter with their left hand to reach for a dish in the bottom rack of a dishwasher to prevent falling. Solving these tasks autonomously is usually done with multi-contact whole-body planners and controllers [bouyarmane2018, padois2017whole].
Figure 1: Talos executes the user’s verbal instructions to (1) lean on a book and (2) reach for an inaccessible cup, yet in its field of view, using our Words2Contact pipeline.
Recent advances in whole-body control using quadratic programming have shown that both torque-controlled robots [henze2016passivity] and position-controlled robots with force/torque sensors [rouxel2024multi] can effectively utilize additional contact points to increase their manipulability and improve their balance, but these control methods require the prior knowledge of the contact locations. This information is usually the output of a contact planning algorithm, where typically a planner decides a sequence of contact locations that enable the robot to solve its task (e.g., walking, manipulating a complex object) [Kumagai2023]. Contacts computation usually relies on visual or 3D perception and environment models to look first for suitable contact surfaces, before deciding whether they are kinematically feasible for the robot. For example, it is common to look for flat areas to place the footsteps in humanoid walking [calvert2022fast].
Unfortunately, selecting contacts, especially when applying forces, often requires an understanding of the world that can hardly be modeled. Some surfaces might be flat, but too fragile for support, like a glass window. Other surfaces might be off-limits for safety reasons, like the wing of an aircraft, or they might be slippery, dirty, or unstable. Overall, in many real-world situations, the choice of support contacts is likely to require human expertise at some point to be deployed outside of a laboratory. Giving the power to human experts to guide the robot and choose the contact locations for them is therefore a very desirable feature.
Human guidance in contact selection is ideal for teleoperated robots in remote maintenance or hazardous scenarios and for collaborative robots cooperating and working side-by-side with humans. For example, a remote operator could instruct the robot to place one end-effector on a wall to lift one foot, and a factory worker could instruct the robot to reach a handle with one end-effector and take a fallen tool with the other one. In these situations, language-based instructions provide a natural communication channel and free the hands of the operator, nor do constrain the human worker to use computer interfaces to instruct the robot on what to do.
Giving instructions in natural language has long been a dream of the robotics community [antoniol1993robust, tellex2020robots, mandery2016using]. For years, this goal eluded researchers due to two main challenges: (1) understanding what a sentence means requires a good intuition of the context and the implicit knowledge, that is, some “common sense” (2); there are countless ways of expressing the same instruction, which prevents the use of simple keywords. To give an illustration, some people might refer to the “Handbook of Robotics” of Fig. 1 as “the big red book”, “the book”,“the book next to your right hand”, “the red thing”, “the big thing in front of you”, and so on.
Large Language Models (LLMs) [zhao2023surveylargelanguagemodels] might be on the verge of solving these challenges for robotics [wang2024largelanguagemodelsrobotics], providing a way to give natural and general instructions to robots. Trained on billions of human-written texts, LLMs exhibit a form of “common sense” that allows them to interpret instructions with their most likely meaning. They are also, perhaps surprisingly, highly versatile, as they are capable of handling instructions or situations not anticipated by the robot designers. Additionally, LLMs naturally process the many ways humans express similar concepts, as they are represented by similar “embeddings”. Visual Language Models (VLMs) are equally appealing, as they can link text to images and vice-versa.
In this paper, we harness the power of Foundation Models (LLMs and VLMs) to instruct a humanoid robot about desired contact locations for increased support in whole-body reaching, an essential skill for solving several downstream tasks.
The robotics community has been working intensely on integrating LLMs with robots since the first demonstrations of ChatGPT (2022). The key challenge is connecting perception, which is continuous, structured, and high-dimensional, to language, which is linear and loosely structured, and then to actions, which are also continuous and depend on the specific robot. While many approaches have been proposed (see Sec. II-B), there is currently no consensus on how to establish this link in the general case.
We present the following key contributions:
  • Words2Contact: a novel pipeline integrating LLMs and VLMs with a multi-contact whole-body controller to identify support contacts from verbal instructions.
  • A benchmark of state-of-the-art LLMs and VLMs for contact prediction.
  • A pilot study showing that users quickly learn to use our system to identify accurate contact locations.
  • Validation of our system on a real Humanoid Robot.
To the best of our knowledge, this paper is the first to address support contact identification from verbal instructions using Foundation Models and demonstrate it with a humanoid robot.
II Related work
II-A Multi-Contact and Whole-Body Control
Multi-contact tasks present both theoretical and practical challenges that must be addressed carefully. In these scenarios, humanoid robots exhibit redundancies in their whole-body kinematic and contact force distribution [seiko], for example a robot can place its hand on a table in multiple positions, with various combinations of force distribution for each configuration. For humanoid robots, maintaining balance and ensuring the feasibility of commanded motions are crucial for safety, and require consideration of the system’s physical constraints. Robustness is essential for real hardware deployment, necessitating real-time control of posture and contact forces.
Low-level control in humanoid robots is often achieved via Quadratic Programming (QP) for whole-body optimization [escande2014hierarchical], allowing fast computation while accounting for constraints and allowing direct control of contact forces on torque-controlled robots [henze2016passivity].
However, such robots are sensitive to model errors, affecting robustness. In this work, we conducted hardware experiments on the full-size humanoid robot Talos [stasse2017talos] controlled in position. In this context, contact forces can be indirectly tracked through admittance schemes [caron2019stair].
Our proposed contact selection method is agnostic to how multi-contact motion is achieved. We utilized our SEIKO (Sequential Equilibrium Inverse Kinematic Optimization) framework [rouxel2024multi] for the multi-contact experiments. SEIKO uses a Sequential QP to formulate a whole-body admittance controller, explicitly modeling joint flexibility to indirectly regulate contact forces on position-controlled robots.
II-B Robotics and Language Models
Prior to the introduction of LLMs, numerous approaches in both language comprehension and generation were explored in robotics [tellex2020robots, cangelosi2010integration]. However, these early methods were limited due to their reliance on rigid, rule-based systems and predefined vocabularies [mandery2016using].
Thanks to their training on a very large dataset, LLMs can answer to a very large set of natural language queries without having been trained on any specific domain.
In particular they can provide high-level plans with some “common sense” by inferring many pieces of context, thus bypassing most of the “frame problem” [dennett1990cognitive]. In robotics, by using well-designed “prompts” that explain the problem to be solved in natural language and the kind of expected output, LLMs were used to find a sequence of pre-learned behaviors [song2023llm, ren2023robotsaskhelpuncertainty], generate Python code to be executed by the robot [singh2022progpromptgeneratingsituatedrobot, liang2023codeaspolicies], or cost functions for a model-based controllers [moweroptimal]. For example, “Inner Monologue” [huang2023reasoninglargelanguagemodels] uses the ability of LLMs to generate task plans and explores embodied reasoning through self-dialogue. “Code-as-Policies” [liang2023codeaspolicies] uses the code generation abilities of LLMs to inform robotic policies directly from natural language descriptions without the need for further training.
In some cases where the desired structure of the output is in a form that LLMs are not inherently able to generate, or if the nature of the problem requires more complex responses, additional fine-tuning may be useful [wu2023embodiedtaskplanninglarge, jin2023alphablockembodiedfinetuningvisionlanguage, kant2022housekeeptidyingvirtualhouseholds]. For example, in “BTGenBot”, behavior trees are generated through LLMs that have been fine-tuned on specialized datasets. The key takeaway is that a well-structured output is beneficial to transition from non-structured high-level instructions to low-level control commands. The drawback of fine-tuning is that it requires large amounts of data, which is time-consuming and resource-intensive to collect, and may introduce biases based on how the data are collected or generated [yu2023largelanguagemodelattributed].
Even though the generated plans are often successful, the ability to use language-based corrections to fix the generated plans generated with minor adjustments during task execution can be very useful. For example, Sharma et al. [sharma2022correctingrobotplansnatural] present a model that integrates natural language and visual feedback to adjust robot planning costs in real-time, enabling more dynamic and responsive adaptation to new tasks. LILAC [cui2023onlinelanguagecorrections] proposes a shared autonomy paradigm that updates the control space in response to continuous user corrections. DROC [zha2024distilling] further advances this paradigm by enabling LLM-based robot policies to respond to, remember, and retrieve feedback efficiently, significantly improving adaptability to natural language instructions. Overall, a correction mechanism that understands general and abstract corrections, such as “a bit to the right”, is essential to ensure the reliability and effectiveness of robotic systems guided by LLMs.
Instead of relying on LLMs solely trained on language, an alternative idea is to use the same learning architectures as LLMs (transformers), but train them on multi-modal robotics data instead of pure text, like in the Robotics Transformers (RT) line of work [brohan2023rt]. A more popular and less compute-intensive approach is to use similar large-scale robotics datasets and incorporate pre-trained language and vision models with a few trained layers to connect the components. OpenVLA [kim2024openvlaopensourcevisionlanguageactionmodel] uses this approach with small open-source models (7 billion parameters compared to GPT-3’s 175 billion.), highlighting the potential for substantial achievements with smaller models.
Regarding humanoid robots, recent research has focused on generating human-like motions from text descriptions, specifically through animation using simulated human-like articulated models [ren2023insactor, motionlcm]. However, for the problem of multi-contact planning, we are only aware of a traditional (pre-LLMs) language processing-based approach, where an n-gram language model is employed. The goal of this model is to learn motion as a sequence of transitions, where each word represents a shape pose, and each sentence represents a motion [mandery2016using].
III Methods
Figure 2: Words2Contact overview: (1) The user provides the first instruction. The Module Selector (Sec. III-B) classifies it as “Prediction”. The Prediction Module (Sec. 4) integrates the user input and the robot’s RGB data to predict a new contact point. (2) The user wants to adjust the predicted contact; their instruction is classified as “correction”. The Correction Module (Sec. III-D) adjusts the previous prediction based on the new user input and the RGB data. (3) The user confirms the corrected contact location: the instruction is classified as “Confirmation”. The Control Module (Sec. III-E) uses the PointCloud, the initial user prompt and the desired contact location to compute the desired 3D contact task, later executed by the SEIKO Multi-Contact whole-body controller.
The Words2Contact pipeline (Fig. 2) unfolds as follows: visual feedback is streamed to the user, who starts by instructing the robot to place a contact in a specified location. The initial prediction resulting from this instruction is displayed to the user. If dissatisfied, the user can either correct the prediction or provide a different instruction until they confirm satisfaction with the updated predicted target. Once the contact location is confirmed, the robot proceeds to execute the contact placement at the specified point using the SEIKO controller.
To achieve this, we split the contact prediction task into three sub-modules, each responsible for a specific sub-task: Prediction, Correction, and Confirmation. This split is crucial for ensuring that even small models will be able to effectively handle each stage of the pipeline.
III-A Prompting LLMs
We use a single LLM and dynamically adjust the system prompt (Fig. 3) at each step of the pipeline. Furthermore, outputs from the LLM are constrained to JSON format to ensure a desired structure that simplifies information extraction, in the open source models we enforce this constraint with grammar-based token sampling and acceptance [willard2023efficient], whereas for the proprietary model we follow the documentation instructions [^1] OpenAI Docs: https://tinyurl.com/openaijson.
For readers unfamiliar with LLMs, it is important to differentiate between the system prompt and the user prompt. The system prompt is an instruction that guides the LLM’s responses, setting the tone, context, and boundaries for the conversation. Each module has a specific system prompt tailored to its task, as illustrated in Fig. 3. In contrast, the user prompt is the natural language input or query provided by the user.
Figure 3: Some examples of system prompts that are utilized by the LLM in our modules. “+5 examples” refers to the 5 examples that are added to the system prompt, as part of the few-shot prompting technique. All the system prompts are available at: %****␣main.tex␣Line␣525␣****https://hucebot.github.io/words2contact_website/.
III-B Module Selector
The Module Selector (Fig. 2) interprets the user’s natural language prompt and classifies it into one of three categories: Prediction, Correction, or Confirmation.
This classification is achieved by combining two key techniques: few-shot prompting [brown2020language] and logits bias [^2] OpenAI Article: https://tinyurl.com/openailogitbias. Few-shot prompting involves providing a system prompt that describes the task that the LLM has to perform, accompanied by examples to guide it in classifying new inputs correctly. For this, and all the following modules, we use $5$-shot prompting, i.e., we provide five examples. Logits bias is a technique used to adjust the output probabilities of logits, specifically for terms such as ‘Prediction’, ‘Correction’, and ‘Confirmation’. This adjustment aims to prioritize the correct classification of inputs into one of these three categories.
III-C Prediction Module
Figure 4: The Prediction Module (Sec. 4): the Prompt-Analyzer is an LLM that analyzes the user’s prompt and returns a JSON file with the chain of thought, list of objects, and position type (absolute or relative). For absolute positions, a point is extracted via language-grounded segmentation. For relative positions, a language-grounded object detection VLM detects bounding box(es), used by the LLM to predict the contact point.
To interpret the desired location implied by the user, we need to have a system that leverages both natural language instructions and visual state feedback. The Prediction module (Fig. 4) combines both Vision Language (VLMs) and a Large Language Model (LLM). We assume that there are two cases of positions that the user might refer to:
  1. 1.
    Absolute Positions: For prompts specifying a contact that is on an object (e.g., “place your hand on the book”).
  2. 2.
    Relative Positions: For prompts where the contact is expressed in terms of its spatial relation to the object(s) (e.g., “left from the box”, “between the cup and the bowl”).
The Prompt Analyzer is responsible for (a) identifying which of the two scenarios the prompt is relevant to and (b) isolating the object’s descriptions mentioned in the prompt so that they can be passed to the VLMs. For complex tasks that involve common sense and math reasoning, chain-of-thought prompting, where the LLM is asked to provide its thought process before reaching a conclusion, has proven to be beneficial [wei2023chainofthought]. We combine few-shot prompting and chain-of-thought reasoning, to achieve better results (see the prompt on Fig. 3-a).
In the case of Absolute Positions (Case 1 in Fig. 4), we use the capability of language-grounded segmentation models to segment images based on natural language descriptions. This allows the system to detect the object regardless of how they are referenced by the user, overcoming the limitations of classic pre-trained segmentation models that either retrieve a mask based on a pre-trained set of labels or return a segmentation of an image without any labeling. CLIPSeg [luddecke2022image], for example, addresses this problem by extending a CLIP model [radford2021learning] with a transformer-based decoder. Once we obtain the segmentation heatmap for the requested object, we determine the coordinates of the contact point in image space using the following metric: $[i_{\text{max}},j_{\text{max}}]=\arg\max_{i,j}\,\mathbf{H}_{ij}$, where $\mathbf{H}$ is the heatmap produced by the language-grounded segmentation model, up-scaled to the size of the original image. While a more sophisticated point sampling technique could be chosen to ensure sufficient space coverage, such considerations are beyond the scope of this work.
In the case of Relative Positions (Case 2 in Fig. 4), we utilize spatial relationships derived from the visual scene and the verbal instruction to determine the contact location. To achieve this, following the same intuition as in the first case, we extract the bounding box(es) using pre-trained open-set object detection [liu2023grounding]. Similarly to grounded segmentation, these models are trained using bounding box annotations and aim at detecting arbitrary classes with the help of language generalization. The representation of a bounding box using natural language is straightforward and thus motivates our approach. For instance, in case 2 of Fig. 4, after we receive a bounding box for the cup, we build the following prompt: “Cup is at [100,150] with width=120 and height=90. Place your hand left from the cup.” The system prompt (Fig. 3-b) contains some basic information about the representation we are following. Additionally, we provide a few examples to ensure that the LLM will accurately interpret the spatial instruction accurately, and will calculate the final contact position successfully. Furthermore, instead of directly outputting a numerical value, the LLM outputs a mathematical expression which is then parsed and calculated using a Python parser. This choice was made because, in preliminary experiments, we noticed a performance increase of around $10\%$ when using this approach instead of having the LLM perform the computation on its own.
III-D Correction Module
When dealing with humanoid robots and contacts, precision is of high importance. The Correction module (Fig. 5) enhances the Words2Contact pipeline by allowing for both minor and major corrections. Similarly to the second scenario of the Prediction Module, we detect the object(s) stated in the user prompt, and then we retrieve their bounding boxes. The main difference is that in the system prompt we mention that the goal now is to correct a user-given position, and in the final user prompt, we include the current target position. Another important point to specify is stating a correction that includes an object (e.g., “Move closer to the cup.”) is optional, and we even support prompts of the form: “Move the target a bit to the right.”. Finally, to provide a more natural interaction with the correction module, we include the conversation history, which allows the operator to make corrections relevant to the previous ones, for example “Move to the right.”, “Now, move twice as much as before.”.
Figure 5: In the Correction Module (Sec. III-D) the LLM detects object descriptions in the user prompt, the VLM identifies their bounding boxes, and then uses them along with the current target position, interaction history, and the user’s instruction to determine a new candidate contact $[i,j]$.
III-E Control Module
Once the user confirms their satisfaction with the displayed contact point, we query the LLM one final time using the end-effector selector prompt (Fig. 3-c), to determine the robot’s end-effector (e.g., right or left hand) and the task type (e.g., support contact or reaching). We extract the 3D position, in camera frame $\mathbf{p}^{cam}=[x,y,z]^{cam}$, from the point cloud and we transform it to robot’s world frame using forward kinematics. A spline-based Cartesian trajectory, starting from the current position $\mathbf{x}^{EE}_{t}$ at time $t$, brings the end-effector EE to the desired contact or reaching point $\mathbf{p}^{rob}$.
The Control Module utilizes SEIKO Retargeting [seiko, seiko2] and Controller [rouxel2024multi] to track the effector position, smoothly establishing new contacts by redistributing contact forces across the effector to participate in the robot’s balance. The SEIKO pipeline takes as input the instantaneous effector target position sampled from the spline, along with the contact force measurements from force-torque sensors, and produces joint positions that are sent to the robot’s hardware. SEIKO is implemented as a model-based Sequential Quadratic Programming (SQP) optimization method to initially compute the desired whole-body configuration based on the target effector position. The Controller then regulates contact forces to achieve multi-contact motion while ensuring robustness. For safety, SEIKO enforces motion feasibility and imposes constraints on joint position and velocity limits, quasi-static balance, and conditions to prevent slipping and tipping. If the pipeline receives an infeasible or unreachable position, the robot will attempt to approach the target as closely as possible without compromising its balance.
Figure 6: In the Control Module (Sec. III-E), the SEIKO Controller commands the robot to realize the desired task with the selected end-effector at the desired contact position.
IV Experiments & Results
IV-A Evaluation of contact prediction using pre-trained models
In this experiment, we benchmark several state-of-the-art pre-trained models (VLMs and LLMs): the goal is to select the best combination for our pipeline, evaluating the impact of the type of model and its size (i.e., the number of parameters) on the prediction performance (mapping user inputs to pixels).
Figure 7: Selected records from our dataset. The yellow masks indicate acceptable contact areas for each given prompt.
To this purpose, we created a new dataset [^3] The dataset can be downloaded from our website: https://hucebot.github.io/words2contact_website/, with $78$ tuples of images (1280$\times$720 pixels), prompts, and manually annotated masks corresponding to the area that satisfies the described contact. The dataset has both indoor and outdoor images, contains different ways of requesting contacts (e.g., lean, place), different end-effectors (e.g., hand, foot), and an even distribution of relative and absolute positions (Fig. 7).
To evaluate the performance of the prediction module and assess the impact of the LLM size, and the VLM choice on the performance of our pipeline, we evaluate the success rate of each combination of several models: For the LLMs, we test Calme-7b-Instruct, mixtao-7bx2-moe and GPT-3.5-turbo. While for the VLM segmentation we chose CLIPSeg [luddecke2022image] and CLIP Surgery [li2023clip], and for the object detection GroundingDINO [liu2023grounding] and Florence-2 [xiao2023florence2]. Random selection was also used to establish a baseline.
The results (Tab. I), show that all the combinations outperform the random point sampling, and that the best combination of models (gpt3.5+GroundingDino+ClipSeg) selects appropriate contacts in about 70% of the absolute cases and 50% of the relative cases. One key finding is that breaking down larger tasks into smaller subtasks enables smaller models to achieve a success rate comparable to larger models, despite their significantly reduced size. This suggests that task structuring can be a valuable strategy in optimizing model performance, especially when computational resources are limited.
TABLE I: Success rate of each combination of foundation models.
Combination Success Rate
LLM VLM ObjectDetection VLM Segmentation absolute relative overall
median [25%, 75%] median [25%, 75%] median [25%, 75%]
Calme-7b-Instruct Florence-2 CLIPSeg $0.67$ [$0.66$, $0.68$] $0.39$ [$0.39$, $0.4$] $0.54$ [$0.39$, $0.66$]
Calme-7b-Instruct Florence-2 CLIP Surgery $0.43$ [$0.42$, $0.45$] $0.42$ [$0.41$, $0.43$] $0.42$ [$0.42$, $0.45$]
Calme-7b-Instruct GroundingDINO CLIPSeg $0.71$ [$0.7$, $0.71$] $0.46$ [$0.45$, $0.48$] $0.59$ [$0.47$, $0.71$]
Calme-7b-Instruct GroundingDINO CLIP Surgery $0.45$ [$0.43$, $0.45$] $0.46$ [$0.43$, $0.48$] $0.45$ [$0.43$, $0.47$]
mixtao-7bx2-moe Florence-2 CLIPSeg $0.66$ [$0.63$, $0.69$] $0.34$ [$0.34$, $0.36$] $0.51$ [$0.34$, $0.64$]
mixtao-7bx2-moe Florence-2 CLIP Surgery $0.43$ [$0.42$, $0.45$] $0.36$ [$0.33$, $0.39$] $0.42$ [$0.36$, $0.45$]
mixtao-7bx2-moe GroundingDINO CLIPSeg $0.66$ [$0.65$, $0.68$] $0.41$ [$0.39$, $0.42$] $0.53$ [$0.41$, $0.66$]
mixtao-7bx2-moe GroundingDINO CLIP Surgery $0.42$ [$0.41$, $0.43$] $0.45$ [$0.41$, $0.47$] $0.42$ [$0.41$, $0.45$]
gpt-3.5 Florence-2 CLIPSeg $0.74$ [$0.72$, $0.74$] $0.39$ [$0.38$, $0.39$] $0.55$ [$0.39$, $0.73$]
gpt-3.5 Florence-2 CLIP Surgery $0.42$ [$0.41$, $0.45$] $0.34$ [$0.34$, $0.39$] $0.41$ [$0.36$, $0.44$]
gpt-3.5 GroundingDINO CLIPSeg $0.71$ [$0.68$, $0.74$] $0.5$ [$0.5$, $0.53$] $\mathbf{0.61}$ [$0.51$, $0.7$]
gpt-3.5 GroundingDINO CLIP Surgery $0.45$ [$0.43$, $0.46$] $0.53$ [$0.5$, $0.54$] $0.47$ [$0.45$, $0.51$]
Results using Random Point Sampling $0.12$ [$0.08$, $0.14$] $0.17$ [$0.11$, $0.21$] $0.13$ [$0.1$, $0.18$]
IV-B Pilot Study - Evaluation of the correction mechanism and usability of the pipeline
Figure 8: Results of the pilot study: all the participants were able to bring the predicted contact close to the target in a few iterations with our system, exhibiting quick learning, both with and without prompt expert guidance. The correction mechanism is very effective in achieving accuracy in contact placement, which is critical for real robot applications.
To evaluate the performance and usability of our system, we conducted a pilot study with 11 volunteer participants ($9$ male, $2$ female, aged $26.27\pm 1.8$ y.o., min $24$, max $30$). All participants had no prior experience with our system. The pilot study was structured in two phases.
In the first phase, participants were presented with a set of $10$ images, randomly sampled from the same dataset as in Sec. IV-A. Each image displayed a random target marked with a circle of 18 pixels radius; the participant’s task was to tell the system how to accurately place a point, marked with another circle of 5 pixels radius, on the designated target using a maximum of 5 prompt steps. Each participant received minimal instructions and, notably, was not informed about the existence of the correction module.
In the second phase, we provided the same participants with an explanation of the system’s functionalities, including two examples of prediction sentences and two correction sentences. With this new knowledge, they instructed the system to identify $10$ targets on $10$ different images.
We measured the distance between the predicted point and the target across prompts, as the distance between the centers of the two circles. The results (Fig. 8) demonstrate a significant improvement in task performance in both phases: all the users quickly learned how to use the system, and were able to bring the point close to the target with few corrections. Their performance was, as expected, better after the prompt expert suggestions (the median distance at the 5th prompt is 21 pixels, which is comparable with the target circle radius).
The quick and accurate target reaching demonstrates the effectiveness of the prediction-correction mechanism in Words2Contact for precise contact identification, essential for real robot applications. Participants also reported high engagement and satisfaction with the system.
IV-C Real Robot experiment
We evaluated Words2Contact with the Talos humanoid robot [stasse2017talos] in four distinct whole-body reaching settings (Fig. 9), with and without corrections, with the following user prompts p:
  • (a)
    1
    Place your right hand on top of the book2with your left hand, reach for the cup”.
  • (b)
    1
    Using your right hand, lean on top of the white surface2reach for the red plate, with the left hand”.
  • (c)
    1
    Place your right hand right from the thing with the wooden handle” (this is a mallet, but the user might not know the name); operator’s correction to avoid a collision: 2Move more to the right”; 3Reach for the nail box, with your left hand”.
  • (d)
    1
    Place your right hand on the white cloth”; 6 corrections 2- 7 guide the target, as the setting lacks distinct objects for relative positioning; 8with the left hand, reach the cheez it box”.
We highlight that the operator completed all the tasks using paraphrases to describe objects (“thing with a wooden handle”) and rare names (“cheez it box”). The robot always performed the task successfully.
Video/Code/Dataset: The video of the robot experiments, the dataset of Sec. IV-A and the software to reproduce Words2Contact are available at https://hucebot.github.io/words2contact_website/.
Figure 9: Talos receives instructions in natural language from an operator. The images show examples of sequences of reaching actions with a support contact in different scenarios: a) book on table, b) dishwasher bottom rack, c) box on table, d) cloth over the fridge. Orange dashed lines indicate the motion of the robot’s end-effector; blue targets indicate the prediction module’s estimated contact locations; yellow targets indicate the corrected target if adjustments were made. Between each execution, the operator confirmed their satisfaction with the predicted target location, but the confirmation steps are excluded for clarity. A video showing the experiments is available at https://hucebot.github.io/words2contact_website/.
V Conclusions & Future Work
Words2Contact is especially useful in scenarios where users cannot directly control the robot, such as in human-robot collaboration settings and teleoperation. For instance, in a collaborative task, the robot could autonomously choose and adjust its hand placement on a shared work surface while the human partner focuses on a different aspect of the task. While the system adapts well in these contexts, the confirmation of the final contact point remains an open question. In this work, we mainly focused on language-based teleoperation, where the operator is remote from the scene and can confirm the location of the final contact point through visual feedback; however an equivalent confirmation system will also be necessary in human-robot collaboration. In our pilot study, users quickly adapted to the system, and real-world experiments demonstrated that Words2Contact delivers satisfactory contact placements, even in challenging environments, through its iterative correction mechanism. Based on these findings, we expect similar performance in human-robot collaboration, where a reliable confirmation system could ensure precise contact point adjustments during shared tasks.
For future work, we aim to reduce user reliance on visual feedback by enabling online corrections and scene-grounded trajectory generation for the end-effectors [shi2024yell, cui2023onlinelanguagecorrections, kwon2024language]. Additionally, improving the system’s performance will involve ensuring motion feasibility through real-time evaluation of inverse kinematics and surface safety using VLMs. For example, predicted contact points will be constrained to areas that are both a) reachable and b) safe, such as avoiding fragile surfaces like glass.
Finally, Words2Contact opens up opportunities for shared or full autonomy by capturing human-robot interaction data during teleoperation to build a robust dataset. This data could be used to fine-tune LLMs or train imitation learning approaches for shared autonomy [rouxel2024flowmatchingimitationlearning], allowing the system to autonomously manage routine tasks while requiring human intervention only for more complex decisions.
[…]